According to company reports, the current mix of colors for M&M dark chocolate candies is: 20% yellow, 10% red, 15% orange, 10% blue, 25% green, with the rest being brown.
A. If a single piece is taken from a bag of 600 pieces, find the following probabilities:
a. P (red)
b. P (brown)
c. P (blue or orange)
d. P (NOT red)
e. P (NOT brown)
f. P (red and green)


B. If two pieces are taken from a 600 piece bag, find the following probabilities:
a. P (both are blue)
b. P (the first is orange)
c. P (first is red and the second is green)
C. How would the process in problem #1A change if we are sampling from a “fun-size” bag which only has 12 pieces?
D. How would the process in problem #1B change if we are sampling from a “fun-size” bag which only has 12 pieces?

Answer by edjones(8007)

100-20-10-15-10-25=20% are brown
600*.2=120 yellow
600*.1= 60 red
600*.15=90 orange
600*.10=60 blue
600*.25=150 green
600*.2=120 brown
.
B)
1) 60/600 * 59/599 = .0098
2) 90/600 = .15
3) 60/600 * 150/599 = .0250
.
Ed